### v1

@@ Abstract
The common observation of a ‘$1/f^{\alpha}$’ power spectrum with $\alpha<2$ constitutes one of the enduring mysteries of condensed matter physics. Here it is shown that a $1/f^{\alpha}$ power spectrum, with $\alpha=3/2$, can arise when an ensemble of two–level systems is coupled to a heat bath by means of a system of Bosonic quasiparticles. The model considered is the classic model of Faughnan and Strandberg of the phonon bottleneck, and the anomalous relaxation is associated with an approximate non-equilibrium steady state of the phonons maintained by slow spin relaxation. It is shown that a frequency-dependent susceptibility can be defined in the steady state and that the prediction $\alpha=3/2$, or the equivalent stretched exponential relaxation with exponent $\beta=2-\alpha=1/2$, is consistent with existing experimental data. These results give an insight into the origins of anomalous relaxation in condensed matter and the practical behaviour of two-level systems.
@@
Figure 1: Exact power spectrum (blue line) compared with Eq. 1 (magenta dashed line). Here $\sigma=10^{5}$ and a power law region with $\alpha=3/2$ extends over $1/\sigma<\omega<\sigma$. The two curves are almost coincident on a log–log scale, illustrating a rare instance where a ‘$1/f^{\alpha}$’ power spectrum can be calculated analytically (here $\alpha=3/2$ and $f=\omega/2\pi$).
@@
These ideas are born out by numerical solution of the coupled equations. Using the example of $a=10^{2},\sigma=10^{5}$, Fig. 2 (upper) illustrates the solution for $v(t)$ compared to $1-\sqrt{2t/\sigma}$ and to $e^{-\sqrt{2t/\sigma}}$, which are very close approximations at small times, and the (exact) inverse Fourier transform [15] of the right hand side of Eq. 1, which captures the whole curve. Fig. 2 (lower) is the corresponding numerical Fourier transform on a log-log scale, confirming a slope very close to 3/23/23 / 2. Fig. 3 illustrates the corresponding $u(t)$ compared to $\frac{d}{dt}e^{-\sqrt{2t/\sigma}}$ (see Eqs. 7 and 8) which is a close approximation at short time after transient effects have died out.
Figure 2: Relaxation of the magnetization difference: numerical solution to Eqs. 2, 3, for $a=10^{2},\sigma=10^{5}$. (a) the function $v(t)$ (blue line) compared to $1-\sqrt{2t/\sigma}$ (gray dotted line), $e^{-\sqrt{2t/\sigma}}$ (gray dashed line) and to the (exact) inverse Fourier transform of the right hand side of Eq. 1 (magenta dashed line). (b) Fast Fourier Transform of $v(t)$ (blue points) compared to a line of slope 3/23/23 / 2 (orange line) on the log-log graph.
@@
for large $\omega<\sigma$ (cf Eq. 1, Eq. 9). This justifies the close approximation of the exact form by Eq. 1, as displayed in Fig. 1. However, the high order terms in $\omega$ in Eq. 15 ensure that the true asymptotic behavior of $|\chi^{\prime\prime}|/\omega$ as $\omega\rightarrow\infty$ is $1/\omega^{2}$, with a crossover to this form at $\omega\sim\sigma$. Thus the anomalous ‘$\alpha=3/2$’ scaling applies to the range $1/\sigma<\omega<\sigma$ and only the divergence of $\sigma$ would extend its range to infinity. Nevertheless, as illustrated in Fig. 4, for finite $\sigma$, the crossover is particularly sharp and has negligible effect on the power law scaling at smaller $\omega$. Were such behavior encountered in experiment, it would surely be interpreted as an asymptotic $1/\omega^{3/2}$ power law as one would have to access unfeasibly small values of $|\chi^{\prime\prime}|/\omega$ to see the crossover.
@@
The value of $\alpha=3/2$ found here, equivalent to stretched exponential ($e^{-t^{\beta}}$) relaxation with $\beta=2-\alpha=1/2$, may be compared to the values found in other magnetic systems. Ising spin glasses [6] display values of $\beta\approx 0.25-0.3$. In single molecule magnets, a value of $\beta=1/2$ has been observed at millikelvin temperatures at resonant tunneling points and interpreted in terms of fluctuating random fields [8, 9]. In spin ice, $\alpha$ evolves from 222 down to about 1.5 at subkelvin temperatures [11], where the value of roughly 1.5 has been associated with the diffusion of emergent magnetic monopoles in a spatially fractal environment [10, 20]. It is perhaps unsurprising that that the few-body systems (paramagnets, single molecule magnets) seem to exhibit exponents that are rational fractions, while the many body systems (spin glasses, spin ice) exhibit more general exponent values.

### v2

@@ Abstract
The common observation of a $1/f^{\alpha}$’ power spectrum with $\alpha<2$ constitutes one of the enduring mysteries of condensed matter physics. Here it is shown that a $1/f^{\alpha}$ power spectrum, with $\alpha=3/2$, can arise when an ensemble of two–level systems is coupled to a heat bath by means of a system of Bosonic quasiparticles. The model considered is the classic model of Faughnan and Strandberg of the phonon bottleneck, and the anomalous relaxation is associated with an approximate non-equilibrium steady state of the phonons maintained by slow spin relaxation. It is shown that a frequency-dependent susceptibility can be defined in the steady state and that the prediction $\alpha=3/2$, or the equivalent stretched exponential relaxation with exponent $\beta=2-\alpha=1/2$, is consistent with existing experimental data. These results give an insight into the origins of anomalous relaxation in condensed matter and the practical behaviour of two-level systems.
@@
Figure 1: Exact power spectrum (blue line) compared with Eq. 1 (magenta dashed line). Here $\sigma=10^{5}$ and a power law region with $\alpha=3/2$ extends over $1/\sigma<\omega<\sigma$. The two curves are almost coincident on a log–log scale, illustrating a rare instance where a $1/f^{\alpha}$’ power spectrum can be calculated analytically (here $\alpha=3/2$ and $f=\omega/2\pi$).
@@
These ideas are born out by numerical solution of the coupled equations. Using the example of $a=10^{2},\sigma=10^{5}$, Fig. 2 (upper) illustrates the solution for $v(t)$ compared to $1-\sqrt{2t/\sigma}$ and to $e^{-\sqrt{2t/\sigma}}$, which are very close approximations at small times, and the (exact) inverse Fourier transform [15] of the right hand side of Eq. 1, which captures the whole curve. Fig. 2 (lower) is the corresponding numerical Fourier transform on a log-log scale, confirming a slope very close to $3/2$. Fig. 3 illustrates the corresponding $u(t)$ compared to $\frac{d}{dt}e^{-\sqrt{2t/\sigma}}$ (see Eqs. 7 and 8) which is a close approximation at short time after transient effects have died out.
Figure 2: Relaxation of the magnetization difference: numerical solution to Eqs. 2, 3, for $a=10^{2},\sigma=10^{5}$. (a) the function $v(t)$ (blue line) compared to $1-\sqrt{2t/\sigma}$ (gray dotted line), $e^{-\sqrt{2t/\sigma}}$ (gray dashed line) and to the (exact) inverse Fourier transform of the right hand side of Eq. 1 (magenta dashed line). (b) Fast Fourier Transform of $v(t)$ (blue points) compared to a line of slope $3/2$ (orange line) on the log-log graph.
@@
for large $\omega<\sigma$ (cf Eq. 1, Eq. 9). This justifies the close approximation of the exact form by Eq. 1, as displayed in Fig. 1. However, the high order terms in $\omega$ in Eq. 15 ensure that the true asymptotic behavior of $|\chi^{\prime\prime}|/\omega$ as $\omega\rightarrow\infty$ is $1/\omega^{2}$, with a crossover to this form at $\omega\sim\sigma$. Thus the anomalous $\alpha=3/2$’ scaling applies to the range $1/\sigma<\omega<\sigma$ and only the divergence of $\sigma$ would extend its range to infinity. Nevertheless, as illustrated in Fig. 4, for finite $\sigma$, the crossover is particularly sharp and has negligible effect on the power law scaling at smaller $\omega$. Were such behavior encountered in experiment, it would surely be interpreted as an asymptotic $1/\omega^{3/2}$ power law as one would have to access unfeasibly small values of $|\chi^{\prime\prime}|/\omega$ to see the crossover.
@@
The value of $\alpha=3/2$ found here, equivalent to stretched exponential ($e^{-t^{\beta}}$) relaxation with $\beta=2-\alpha=1/2$, may be compared to the values found in other magnetic systems. Ising spin glasses [6] display values of $\beta\approx 0.25-0.3$. In single molecule magnets, a value of $\beta=1/2$ has been observed at millikelvin temperatures at resonant tunneling points and interpreted in terms of fluctuating random fields [8, 9]. In spin ice, $\alpha$ evolves from $2$ down to about 1.5 at subkelvin temperatures [11], where the value of roughly 1.5 has been associated with the diffusion of emergent magnetic monopoles in a spatially fractal environment [10, 20]. It is perhaps unsurprising that that the few-body systems (paramagnets, single molecule magnets) seem to exhibit exponents that are rational fractions, while the many body systems (spin glasses, spin ice) exhibit more general exponent values.